ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 58E09, 76D15 Loginov, Boris V., Ulyanovsk TU, Russia Bifurcation theory methods in the problem about capillarygravity waves in spatial layer of floating fluid Bifurcation theory under group invariance conditions is applied to the free boundary value problem for the system of differential equations describing the surface waves in floating fluid. The linearized operator is sought by the change of variables “straightening” unknown free boundary. Bifurcating solutions asymptotics are calculated and their stability is investigated. Reference: B. V. Loginov, Group analysis methods for construction and investigation of the BEq, Appl. of Math. 37 (1992),4,241-248.